Find the 95% confidence interval for the variance and standard deviation for the sugar content in ice cream (in mg) if a sample of nine servings has a variance of 36.
1
Expert's answer
2021-10-14T18:31:26-0400

The critical values for "\alpha = 0.05" and "df =n-1=9-1= 8" degrees of freedom are "\chi^2_L=\chi^2_{1-\alpha/2,n-1}=2.1797" and "\chi^2_L=\chi^2_{\alpha/2,n-1}=17.5345."

The corresponding confidence interval is computed as shown below:


"CI(Variance)=\bigg(\dfrac{(n-1)s^2}{\chi^2_{\alpha/2,n-1}},\dfrac{(n-1)s^2}{\chi^2_{1-\alpha/2,n-1}}\bigg)"

"=\bigg(\dfrac{8(36)}{17.5345},\dfrac{8(36)}{2.1797}\bigg)"

"=(16.4248, 132.1283)"

Now that we have the limits for the confidence interval, the limits for the 95% confidence interval for the population standard deviation are obtained by simply taking the squared root of the limits of the confidence interval for the variance, so then:


"CI(Standard\ Deviation)=\bigg(\sqrt{\dfrac{288}{17.5345}}, \sqrt{\dfrac{288}{2.1797}}\bigg)"

"=(4.0527,11.4947)"

Therefore, based on the data provided, the 95% confidence interval for the population variance is "16.4248 < \sigma^2 < 132.1283," and the 95% confidence interval for the population standard deviation is "4.0527 < \sigma < 11.4947."